OpenStudy (sparklestaraa):

In a certain town, the inches of rainfall, x, on a certain day is described by two statements: If 3 times the rainfall is increased by 2, the rainfall is still less than 26 inches. 5 plus twice the rainfall is greater than 11 inches. Part A: Create a compound inequality to represent the rainfall range. Part B: Can the rainfall in this town be 4 inches? Justify your answer by solving the inequalities in Part A. Part C: The average rainfall in another town is 5 inches, but the actual rainfall is within 4 inches of the average. Write and solve an inequality to find the r

OpenStudy (sparklestaraa):

@dtan5457

OpenStudy (danjs):

three times the rainfall is increased by 2 3*x + 2

OpenStudy (danjs):

It is still less than 26 inches 3*x + 2 < 26

OpenStudy (danjs):

Five plus twice the rainfall 5 + 2*x

OpenStudy (danjs):

Is greater than 11 inches 5 + 2*x > 11

OpenStudy (danjs):

Need to solve each of those for x, the amount of rainfall first

OpenStudy (danjs):

3x + 2< 26 5 + 2x > 11

OpenStudy (dtan5457):

lol you have @DanJS helping you. no need for me ;p

OpenStudy (danjs):

3x < 24 x < 8 2x >6 x > 3

OpenStudy (danjs):

so the amount of rainfall is bigger than 3 but less than 8, in other words 3<x<8

OpenStudy (danjs):

you see what i did?

OpenStudy (sparklestaraa):

wait xD I'm still reading and understanding

OpenStudy (danjs):

Take it one sentence at a time

OpenStudy (sparklestaraa):

3x <24 where is that from??

OpenStudy (danjs):

Do you understand how the 2 inequalities were made from the problem paragraph first?

OpenStudy (danjs):

3x + 2< 26 5 + 2x > 11

OpenStudy (sparklestaraa):

ohh yes because what it said signaled different signs etc.

OpenStudy (danjs):

u see how those 2 inequalities were made?

OpenStudy (sparklestaraa):

yes :D

OpenStudy (sparklestaraa):

i understand how u made these 3x + 2< 26 5 + 2x > 11

OpenStudy (danjs):

ok, you have to solve each of them for x, from the first one 3x + 2 < 26 (take 2 away from both sides) 3x <24 (divide by 3) x<8

OpenStudy (sparklestaraa):

thats all in part a still ryt xD

OpenStudy (danjs):

oh Part A is just writing down the 2 inequalities, Part B is solving like now

OpenStudy (danjs):

the second 5 + 2x >11 take away 5 both sides 2x > 6 x>3

OpenStudy (danjs):

so x is bigger than 3 but less than 8

OpenStudy (danjs):

3 < x < 8

OpenStudy (sparklestaraa):

r we onto part c??

OpenStudy (danjs):

Part B) can the rainfall x be 4 inches?

OpenStudy (danjs):

we just solved and got 3<x<8

OpenStudy (danjs):

the rainfall is between 3 and 8 inches

OpenStudy (sparklestaraa):

ohhhhhh :D

OpenStudy (sparklestaraa):

ok i understand that

OpenStudy (danjs):

Part C) r = rainfall

OpenStudy (danjs):

The average rainfall in another town is 5 inches, but the actual rainfall is within 4 inches of the average. Write and solve an inequality to find the r

OpenStudy (danjs):

The actual rainfall r, is within 4 inches of the average. average is 5 , 4 inches in either direction of that is (5-4) and (5+4) the actual rainfall is between (5-4) < r < (5+4) 1 < r < 9

OpenStudy (sparklestaraa):

i wrote this Part A - 3*x+2 3*x+2 <26 5+2*x 5+2*x > 11 3x+2<26 5+2x>11 Part B- 3x+2<26 3x <24 x<8 5+2x>11 2x>6 x>3 3<x<8 The rainfall is between 3 and 8 inches. Part C- R=rainfall The actual rainfall (r) is within 4 inches of the average. average=5, 4 inches in each direction is (5-4) & (5+4) Actual rainfall is between (5-4) < r < (5+4) 1<r<9 Right?

OpenStudy (sparklestaraa):

@DanJS

OpenStudy (danjs):

looks good, you got it

OpenStudy (danjs):

Just say for part B, that the rainfall can be 4 inches because it has to be between 3 and 8 inches from the equations

OpenStudy (sparklestaraa):

so don't write any problem just say that???

OpenStudy (danjs):

no still write everything you did, just include that last bit i wrote

OpenStudy (danjs):

It askes if the rain can be 4, you didnt say yes or no

OpenStudy (sparklestaraa):

ohhh ok LD

OpenStudy (sparklestaraa):

:D

OpenStudy (sparklestaraa):

thanksss ok can u check this one

OpenStudy (sparklestaraa):

Part A: Solve A = 9/2 (x + 28) for x. Part B: Determine the value of x when A = 135. Part C: Solve -np - 40 > 10 for n. Show your work.

OpenStudy (sparklestaraa):

and i got- Part A- A=9x/2 + 9*28/2 A= 9x/2 +126 Part B- 135 = 9x/2 +126 135-126=9x/2 9=9x/2 x=2 Part C- -np-40>10 -np >10+40 -np > 50 -n >50/p n < -50/p @DanJS

OpenStudy (danjs):

looks good, you got it

OpenStudy (sparklestaraa):

:D thanks do u know anything on linear functions